# thanks: Jong-Won Leethanks: Jung-Woo Lee

### v1

@@ I Introduction
Constraining the symmetry energy term in the nuclear equation of state (EoS) is of central importance in the field of nuclear physics and astrophysics. The symmetry-energy potential affects the binding energy of a system to favor the formation of a more isospin-symmetric nucleus. This behavior opposes the trend of the Coulomb interaction, which favors the formation of neutron-rich nuclei. The interplay of these two potential terms limits the isospin asymmetry, $\delta=(N-Z)/A$, of nuclei. In astrophysics, the EoS serves as input to the Tolman-Oppenheimer-Volkoff (TOV) equation [Oppenheimer:1939ne, Tolman:1939jz] for which the solution determines the structure of static compact astronomical objects. Particularly in neutron stars, the density dependence of the symmetry energy ($\mathcal{E}_{sym}=S(\rho)\delta^{2}$) significantly impacts the pressure-density relation, which determines various bulk properties such as internal structure, mass-radius relations, and cooling mechanisms [Lattimer:2004pg, Lattimer:2016].
@@
Various efforts have been made in astronomical observations [Doroshenko:2022nwp] to constrain the EoS of cold dense matter. In particular, the tidal deformability extracted from the recent gravitational wave observations [LIGOScientific:2017vwq] and the mass-radius correlations from the NICER collaboration [Riley:2019, Miller:2019] have placed important constraints on the total pressure of neutron stars at supra-saturation density. However, large uncertainties for the symmetry energy [Tsang:2020lmb, Tsang:2019mlz] remain near twice saturation density due to the lack of good quality ta [Tsang:2023vhh].
@@
Heavy-ion collisions (HICs) provide a terrestrial means to create nuclear mediums at a wide range of densities. At intermediate energies, particle production in HICs is a consequence of dynamical processes in the early stage ($\sim 10^{-22}$ s) of the reaction when the projectile and target nuclei overlap, followed by statistical decay processes which extend to a much longer time scale. Depending on the incident energy and the impact parameter, the system can be compressed to around twice the saturation density. Violent collisions among the constituent nucleons cause the system to thermalize [Ono:2019jxm]. In this dense, hot collision center, the dynamics of the nucleons and light clusters are significantly affected by the symmetry energy [Chen:2004kj]. Light clusters such as deuterons, tritons, and $\alpha$ particles are continuously formed and disintegrated. Depending on density and momentum, many clusters can co-exist and contribute to the dynamics of the system [Ono:2016xun]. The compressed system then expands gradually until it reaches freeze-out, where particles stop interacting with each other.
@@
Transport theory has been the main model to simulate nucleus-nucleus collisions and extract physics from the dynamical processes [TMEP:2022xjg]. In this framework, the equation of motion of nucleons in a nuclear medium evolves under the effect of a self-consistent mean-field potential and the two-body scattering. Important physics, such as the density and isospin dependence of the symmetry energy, can be inferred by comparing observables from experimental data to transport model predictions with different prescribed mean-field potential parameters. For instance, a double neutron-proton yield ratio has been proposed as a sensitive probe to the density dependence of $S(\rho)$ [Coupland:2014gya, Morfouace:2019jky].
@@
In nuclear collisions below 1 $\mathrm{GeV}/\mathrm{nucleon}$, a significant portion of nucleons are emitted in the form of light clusters, which indicates that cluster correlations should be explicitly considered in the scattering processes. Unlike most transport models where the cluster is only recognized at certain freeze-out configurations, $\mathrm{AMD}$ directly incorporates cluster correlations in the final state of two-nucleon collisions during the dynamical evolution of the system [Ono:2020zsx, Ikeno:2016xpr]. After the freeze-out, the excited fragments will undergo sequential decay, which leads to a final state composed of light particles and fragments [Tan:2003bi]. These de-excited fragments are compared to data. $\mathrm{AMD}$ has successfully reproduced particle spectra [SpiRIT:2021och, SpiRIT:2022sqt, INDRA:2023myq], spectral ratios [Ono:2016xun], charge distributions [Ono:2019jxm, Ono:2020zsx] and stopping observables such as the portion of energy transferred from projectile to target in the transverse direction [Ono:2020zsx] observed in various reactions.
@@
In transport models, the nucleon-nucleon scattering cross-section $\sigma_{\mathrm{NN}}$ accounts for the stochastic nucleon-nucleon collision in a nuclear medium. It is well established that the magnitude of $\sigma_{\mathrm{NN}}$ is reduced compared to that in free space [Danielewicz:2002he, Barker:2016hqv, FOPI:2010xrt, INDRA:2002kji]. However, the reduction is not well constrained, and different prescriptions have been used to describe reactions at different densities and energies [FOPI:2010xrt, SpiRIT:2023htl]. In this work, $\mathrm{AMD}$ calculations with different reductions in $\tilde{\sigma}_{\mathrm{NN}}$ are used to describe transverse momentum spectra of light-charged particles from 40,48Ca + 58,64Ni reaction at incident energies of $56$ and $140$ $\mathrm{MeV}/\mathrm{nucleon}$. A direct comparison between data and model calculations provides information about the dependence of the in-medium effect on reaction conditions. To avoid confusion with the free nucleon-nucleon cross-section, we adopt the symbol $\tilde{\sigma}_{\mathrm{NN}}$ to indicate the modified $\sigma_{\mathrm{NN}}$ used in our work.
@@ II Experiment
The experiment was conducted at the National Superconducting Cyclotron Laboratory (NSCL) at Michigan State University. Primary beams of $^{40,48}\mathrm{Ca}$ at $56$, $140$ $\mathrm{MeV}/\mathrm{nucleon}$ were impinged on targets of 5.0 and 5.3 mg/cm2 of 58,64Ni respectively. Table 1 gives a summary of the characteristics of the studied systems. The detection system consists of three parts: the impact parameter detection system (Microball) [Sarantites:1996], the charged-particle detection system (HiRA10) [Wallace:2007, DellAquila:2019rat], and the neutron detection system with charged particle vetos [zecher:1997, Zhu:2020]. Details of the experiment setup can be found in Ref. [Paneru:2023, Sweany:2020abc, Teh:2023szo]. In this study, we mainly focus on the analysis of charged particles from central collisions detected in HiRA10 in coincidence with particles detected in the Microball.
@@
The impact parameter of each event can be estimated from the multiplicity of particles, $N_{C}$, that hit the Microball. A detailed description of this method can be found in [Sweany:2020abc]. During the experiments, only events with Microball multiplicity greater than 5 were recorded. To select central events with $b\lesssim 3$ fm, a software gate on $N_{C}\geq 10$ for reactions at $56$ $\mathrm{MeV}/\mathrm{nucleon}$ and $N_{C}\geq 14$ at $140$ $\mathrm{MeV}/\mathrm{nucleon}$ was applied in the analysis. The detailed procedure will be presented in Sec. III.3.
@@
In each telescope, a $1.5\,\mathrm{mm}$ thick Double-Sided Silicon Strip Detector (DSSD) serves as the $\Delta E$ detector. Each of the front and back sides of the DSSD is composed of 32 strips with a spacing of $1.95\,\mathrm{mm}$. The strips in the front and back sides are oriented perpendicularly to each other, creating a measurement grid with an angular resolution better than $0.5^{\circ}$ [DellAquila:2019rat]. Followed by the DSSD is a composite of four $10\,\mathrm{cm}$-thick CsI(Tl) crystals (E-detector), a major upgrade from the previous $4\,\mathrm{cm}$ design [Wallace:2007]. The upgrade allows charged particles with $E_{\mathrm{lab}}<100$ $\mathrm{MeV}/\mathrm{nucleon}$ to completely deposit their energies. We identify $p$, $d$, $t$, $^{3}{\mathrm{He}}$ and $\alpha$ particles with high statistics. For simplicity, CsI(Tl) will be referred to as CsI in the rest of this manuscript. Details in the calibrations and performance of HiRA10 can be found in [DellAquila:2019rat, Sweany:2021hjd].
@@ III.1 Antisymmetrized Molecular Dynamics
The Antisymmetrized Molecular Dynamics (AMD) model [Ono:2020zsx, SpiRIT:2020sfn, SpiRIT:2021och, SpiRIT:2022sqt, TMEP:2022xjg] is used to simulate the dynamic phase of the reaction systems up to $300$ $\mathrm{fm}/c$ in reactions at $140$ $\mathrm{MeV}/\mathrm{nucleon}$, and up to $500$ $\mathrm{fm}/c$ in reactions at $56$ $\mathrm{MeV}/\mathrm{nucleon}$. The average emission times for low-momentum $\alpha$ particles with $p_{\mathrm{T}}/\mathrm{A}<200$ $\mathrm{MeV}/c$ are slightly below $200$ and $120$ $\mathrm{fm}/c$ in $^{48}\mathrm{Ca}+{}^{64}\mathrm{Ni}$ at $56$ and $140\,\mathrm{MeV}/\mathrm{nucleon}$, respectively. This justifies the selected time limit when the dynamical phase ends and the excited fragments undergo statistical decay [Puhlhofer:1977zz, Maruyama:1992bh]. The impact parameter ranges from 0 to 88 $\mathrm{fm}$ following a geometric distribution. The SLy4 effective interaction [Chabanat:1997, Chabanat:1997un] was employed in this work. It corresponds to the symmetry energy slope parameter $L=46\,\mathrm{MeV}$ and exhibits a quadratic momentum dependence of the mean field, characterized by an effective mass $m^{*}/m=0.70$ in symmetric matter at saturation density.
The ability to describe cluster formation is one of the distinctive features of $\mathrm{AMD}$ model, compared to other transport model approaches [TMEP:2022xjg, Ono:2019jxm]. Cluster correlation is implemented in the final state of each two-nucleon collision ($N_{1}+N_{2}+B_{1}+B_{2}\rightarrow C_{1}+C_{2}$), where $N_{1}$ and $N_{2}$ are the colliding nucleons. The process includes the special cases where a surrounding particle $B_{1}$ and/or $B_{2}$ might be empty [Ono:2020zsx, Ikeno:2016xpr]. For a specific configuration ($C_{1},C_{2}$), the collision cross section is written as
@@
where $p_{i}$ and $v_{i}$ refer to the initial relative momentum and velocity between the scattered nucleons. The relative momentum vector between the nucleons after momentum transfer is denoted as $(p_{f},\Omega)$, where $p_{f}$ is determined from the energy conservation for the adopted effective interaction. This implies that the velocity after momentum transfer $v_{f}=\partial E/\partial p_{f}$, is also dependent on the effective interaction. Any modification to the momentum-dependence of the mean field in Eq. (1) will be reflected in the factor of $p_{i}/v_{i}$ and $p_{f}/v_{f}$. The overlap probability factor for cluster formation, $P(C_{1},C_{2},p_{f},\Omega)$, is defined by considering the non-orthogonality between the states of different configurations [Ikeno:2016xpr, Ono_2013].
@@
where $\tilde{\sigma}_{\mathrm{NN}}$ is parameterized [Coupland:2011px, Danielewicz:2002he] as
@@
In these formulas, the right-hand side of Eq. (2) is evaluated at an average of $p_{\text{i}}$ and $p_{\text{f}}$, and $\rho^{\prime}$ is a density with a momentum cut as defined by Eq. (161) in Ref. [TMEP:2022xjg]. The free parameter $\eta$ in Eq. (4) can be adjusted to control the degree of reduction of the nucleon-nucleon matrix element. Figure 2 illustrates how varying $\eta$ impacts the $\tilde{\sigma}_{np}$ at saturation density.
@@ III.2 Experimental filter
The sharp cutoff method [Cavata:1990gk] has been widely applied in experiments to select central events from observables such as $N_{C}$ and total kinetic energy, which exhibits a negative correlation with $b$. However, the stochastic nature of nucleon-nucleon collisions implies the existence of large fluctuations in such observables for a given $b$ in the transport model. This poses an ambiguity when comparing data with the prediction of the model. Alternative methods include using machine learning [Tsang:2021rku] to map $b$ from a selection of observables and the model-independent method based on Bayes’ theorem [INDRA:2020kyj, Chen:2022kxe, INDRA:2023myq, INDRA-FAZIA:2024rxc]. In this work, we adopt the simple approach by applying the experiment condition in $\mathrm{AMD}$ to select events in the same manner as in the experiment.
Figure 1 shows the hit pattern of charged particles falling within the coverage of the Microball in polar ($\theta$) and azimuthal ($\phi$) angles in the laboratory frame. Each square represents the approximate coverage of a single CsI crystal in the Microball. The particle hits gradually decrease from forward to backward polar angles. In the experimental setup, the front of each CsI in the Microball was covered with a thin foil of Sn/Pb to block electrons generated in the HICs. Particles without sufficient energy to punch through the foils are not detected. Furthermore, due to the fact that a single CsI crystal can not resolve multiple signals, only a single hit is recorded in each crystal in the same event. Figure 3 shows that the probability of the events with large $b$ drops significantly as $N_{C}$ increases, consistent with [INDRA:2020kyj].
Figure 3: Differential cross section $d\sigma(N_{C})/db$ from AMD calculation. The dashed lines are fits to the distribution with the equation $2\pi b/(1+\exp((b-b_{0})/\Delta b))$, see Ref. [INDRA:2020kyj].
@@ IV Results and discussion
The implementation of in-medium effects is not universal across reactions at different beam energies. In our calculation, $\eta=0.35$ is consistent with data in the reaction at $56$ $\mathrm{MeV}/\mathrm{nucleon}$ and $\eta=0.85$ is consistent with data in the reaction at $140$ $\mathrm{MeV}/\mathrm{nucleon}$, indicating more suppression in the medium in the reaction at $56$ $\mathrm{MeV}/\mathrm{nucleon}$. It implies that the in-medium effect depends not only on the local density but also reflects the dynamic situation of the medium, such as a nontrivial phase space distribution. However, previous study in $\mathrm{Sn}+\mathrm{Sn}$ reactions at $270$ $\mathrm{MeV}/\mathrm{nucleon}$ showed that the $\mathrm{AMD}$ calculation cannot simultaneously reproduce the rapidity distributions and the transverse momentum spectra of light charged particles [SpiRIT:2021och, SpiRIT:2022sqt] by only adjusting parameters of the NN collision matrix element.

### v2

@@ I Introduction
Constraining the symmetry energy term in the nuclear equation of state (EoS) is of central importance in the field of nuclear physics and astrophysics. The symmetry-energy potential affects the binding energy of a system to favor the formation of a more isospin-symmetric nucleus. This behavior opposes the trend of the Coulomb interaction, which favors the formation of neutron-rich nuclei. The interplay of these two potential terms limits the isospin asymmetry, $\delta=(N-Z)/A$, of nuclei. In astrophysics, the EoS serves as input to the Tolman-Oppenheimer-Volkoff (TOV) equation [30, 40] for which the solution determines the structure of static compact astronomical objects. Particularly in neutron stars, the density dependence of the symmetry energy ($\mathcal{E}_{sym}=S(\rho)\delta^{2}$) significantly impacts the pressure-density relation, which determines various bulk properties such as internal structure, mass-radius relations, and cooling mechanisms [20, 21].
@@
Various efforts have been made in astronomical observations [13] to constrain the EoS of cold dense matter. In particular, the tidal deformability extracted from the recent gravitational wave observations [1] and the mass-radius correlations from the NICER collaboration [35, 24] have placed important constraints on the total pressure of neutron stars at supra-saturation density. However, large uncertainties for the symmetry energy [43, 45] remain near twice saturation density due to the lack of good quality ta [44].
@@
Heavy-ion collisions (HICs) provide a terrestrial means to create nuclear mediums at a wide range of densities. At intermediate energies, particle production in HICs is a consequence of dynamical processes in the early stage ($\sim 10^{-22}$ s) of the reaction when the projectile and target nuclei overlap, followed by statistical decay processes which extend to a much longer time scale. Depending on the incident energy and the impact parameter, the system can be compressed to around twice the saturation density. Violent collisions among the constituent nucleons cause the system to thermalize [28]. In this dense, hot collision center, the dynamics of the nucleons and light clusters are significantly affected by the symmetry energy [6]. Light clusters such as deuterons, tritons, and $\alpha$ particles are continuously formed and disintegrated. Depending on density and momentum, many clusters can co-exist and contribute to the dynamics of the system [27]. The compressed system then expands gradually until it reaches freeze-out, where particles stop interacting with each other.
@@
Transport theory has been the main model to simulate nucleus-nucleus collisions and extract physics from the dynamical processes [47]. In this framework, the equation of motion of nucleons in a nuclear medium evolves under the effect of a self-consistent mean-field potential and the two-body scattering. Important physics, such as the density and isospin dependence of the symmetry energy, can be inferred by comparing observables from experimental data to transport model predictions with different prescribed mean-field potential parameters. For instance, a double neutron-proton yield ratio has been proposed as a sensitive probe to the density dependence of $S(\rho)$ [10, 25].
@@
In nuclear collisions below 1 $\mathrm{GeV}/\mathrm{nucleon}$, a significant portion of nucleons are emitted in the form of light clusters, which indicates that cluster correlations should be explicitly considered in the scattering processes. Unlike most transport models where the cluster is only recognized at certain freeze-out configurations, $\mathrm{AMD}$ directly incorporates cluster correlations in the final state of two-nucleon collisions during the dynamical evolution of the system [29, 17]. After the freeze-out, the excited fragments will undergo sequential decay, which leads to a final state composed of light particles and fragments [38]. These de-excited fragments are compared to data. $\mathrm{AMD}$ has successfully reproduced particle spectra [19, 22, 14], spectral ratios [27], charge distributions [28, 29] and stopping observables such as the portion of energy transferred from projectile to target in the transverse direction [29] observed in various reactions.
@@
In transport models, the nucleon-nucleon scattering cross-section $\sigma_{\mathrm{NN}}$ accounts for the stochastic nucleon-nucleon collision in a nuclear medium. It is well established that the magnitude of $\sigma_{\mathrm{NN}}$ is reduced compared to that in free space [11, 2, 34, 16]. However, the reduction is not well constrained, and different prescriptions have been used to describe reactions at different densities and energies [34, 42]. In this work, $\mathrm{AMD}$ calculations with different reductions in $\tilde{\sigma}_{\mathrm{NN}}$ are used to describe transverse momentum spectra of light-charged particles from 40,48Ca + 58,64Ni reaction at incident energies of $56$ and $140$ $\mathrm{MeV}/\mathrm{nucleon}$. A direct comparison between data and model calculations provides information about the dependence of the in-medium effect on reaction conditions. To avoid confusion with the free nucleon-nucleon cross-section, we adopt the symbol $\tilde{\sigma}_{\mathrm{NN}}$ to indicate the modified $\sigma_{\mathrm{NN}}$ used in our work.
@@ II Experiment
The experiment was conducted at the National Superconducting Cyclotron Laboratory (NSCL) at Michigan State University. Primary beams of $^{40,48}\mathrm{Ca}$ at $56$, $140$ $\mathrm{MeV}/\mathrm{nucleon}$ were impinged on targets of 5.0 and 5.3 mg/cm2 of 58,64Ni respectively. Table 1 gives a summary of the characteristics of the studied systems. The detection system consists of three parts: the impact parameter detection system (Microball) [36], the charged-particle detection system (HiRA10) [46, 12], and the neutron detection system with charged particle vetos [48, 49]. Details of the experiment setup can be found in Ref. [31, 33, 39]. In this study, we mainly focus on the analysis of charged particles from central collisions detected in HiRA10 in coincidence with particles detected in the Microball.
@@
The impact parameter of each event can be estimated from the multiplicity of particles, $N_{C}$, that hit the Microball. A detailed description of this method can be found in [33]. During the experiments, only events with Microball multiplicity greater than 5 were recorded. To select central events with $b\lesssim 3$ fm, a software gate on $N_{C}\geq 10$ for reactions at $56$ $\mathrm{MeV}/\mathrm{nucleon}$ and $N_{C}\geq 14$ at $140$ $\mathrm{MeV}/\mathrm{nucleon}$ was applied in the analysis. The detailed procedure will be presented in Sec. III.3.
@@
In each telescope, a $1.5\,\mathrm{mm}$ thick Double-Sided Silicon Strip Detector (DSSD) serves as the $\Delta E$ detector. Each of the front and back sides of the DSSD is composed of 32 strips with a spacing of $1.95\,\mathrm{mm}$. The strips in the front and back sides are oriented perpendicularly to each other, creating a measurement grid with an angular resolution better than $0.5^{\circ}$ [12]. Followed by the DSSD is a composite of four $10\,\mathrm{cm}$-thick CsI(Tl) crystals (E-detector), a major upgrade from the previous $4\,\mathrm{cm}$ design [46]. The upgrade allows charged particles with $E_{\mathrm{lab}}<100$ $\mathrm{MeV}/\mathrm{nucleon}$ to completely deposit their energies. We identify $p$, $d$, $t$, $^{3}{\mathrm{He}}$ and $\alpha$ particles with high statistics. For simplicity, CsI(Tl) will be referred to as CsI in the rest of this manuscript. Details in the calibrations and performance of HiRA10 can be found in [12, 37].
@@ III.1 Antisymmetrized Molecular Dynamics
The Antisymmetrized Molecular Dynamics (AMD) model [29, 18, 19, 22, 47] is used to simulate the dynamic phase of the reaction systems up to $300$ $\mathrm{fm}/c$ in reactions at $140$ $\mathrm{MeV}/\mathrm{nucleon}$, and up to $500$ $\mathrm{fm}/c$ in reactions at $56$ $\mathrm{MeV}/\mathrm{nucleon}$. The average emission times for low-momentum $\alpha$ particles with $p_{\mathrm{T}}/\mathrm{A}<200$ $\mathrm{MeV}/c$ are slightly below $200$ and $120$ $\mathrm{fm}/c$ in $^{48}\mathrm{Ca}+{}^{64}\mathrm{Ni}$ at $56$ and $140\,\mathrm{MeV}/\mathrm{nucleon}$, respectively. This justifies the selected time limit when the dynamical phase ends and the excited fragments undergo statistical decay [32, 23]. The impact parameter ranges from 0 to 88 $\mathrm{fm}$ following a geometric distribution. The SLy4 effective interaction [5, 4] was employed in this work. It corresponds to the symmetry energy slope parameter $L=46\,\mathrm{MeV}$ and exhibits a quadratic momentum dependence of the mean field, characterized by an effective mass $m^{*}/m=0.70$ in symmetric matter at saturation density.
The ability to describe cluster formation is one of the distinctive features of $\mathrm{AMD}$ model, compared to other transport model approaches [47, 28]. Cluster correlation is implemented in the final state of each two-nucleon collision ($N_{1}+N_{2}+B_{1}+B_{2}\rightarrow C_{1}+C_{2}$), where $N_{1}$ and $N_{2}$ are the colliding nucleons. The process includes the special cases where a surrounding particle $B_{1}$ and/or $B_{2}$ might be empty [29, 17]. For a specific configuration ($C_{1},C_{2}$), the collision cross section is written as
@@
where $p_{i}$ and $v_{i}$ refer to the initial relative momentum and velocity between the scattered nucleons. The relative momentum vector between the nucleons after momentum transfer is denoted as $(p_{f},\Omega)$, where $p_{f}$ is determined from the energy conservation for the adopted effective interaction. This implies that the velocity after momentum transfer $v_{f}=\partial E/\partial p_{f}$, is also dependent on the effective interaction. Any modification to the momentum-dependence of the mean field in Eq. (1) will be reflected in the factor of $p_{i}/v_{i}$ and $p_{f}/v_{f}$. The overlap probability factor for cluster formation, $P(C_{1},C_{2},p_{f},\Omega)$, is defined by considering the non-orthogonality between the states of different configurations [17, 26].
@@
where $\tilde{\sigma}_{\mathrm{NN}}$ is parameterized [9, 11] as
@@
In these formulas, the right-hand side of Eq. (2) is evaluated at an average of $p_{\text{i}}$ and $p_{\text{f}}$, and $\rho^{\prime}$ is a density with a momentum cut as defined by Eq. (161) in Ref. [47]. The free parameter $\eta$ in Eq. (4) can be adjusted to control the degree of reduction of the nucleon-nucleon matrix element. Figure 2 illustrates how varying $\eta$ impacts the $\tilde{\sigma}_{np}$ at saturation density.
@@ III.2 Experimental filter
The sharp cutoff method [3] has been widely applied in experiments to select central events from observables such as $N_{C}$ and total kinetic energy, which exhibits a negative correlation with $b$. However, the stochastic nature of nucleon-nucleon collisions implies the existence of large fluctuations in such observables for a given $b$ in the transport model. This poses an ambiguity when comparing data with the prediction of the model. Alternative methods include using machine learning [41] to map $b$ from a selection of observables and the model-independent method based on Bayes’ theorem [15, 7, 14, 8]. In this work, we adopt the simple approach by applying the experiment condition in $\mathrm{AMD}$ to select events in the same manner as in the experiment.
Figure 1 shows the hit pattern of charged particles falling within the coverage of the Microball in polar ($\theta$) and azimuthal ($\phi$) angles in the laboratory frame. Each square represents the approximate coverage of a single CsI crystal in the Microball. The particle hits gradually decrease from forward to backward polar angles. In the experimental setup, the front of each CsI in the Microball was covered with a thin foil of Sn/Pb to block electrons generated in the HICs. Particles without sufficient energy to punch through the foils are not detected. Furthermore, due to the fact that a single CsI crystal can not resolve multiple signals, only a single hit is recorded in each crystal in the same event. Figure 3 shows that the probability of the events with large $b$ drops significantly as $N_{C}$ increases, consistent with [15].
Figure 3: Differential cross section $d\sigma(N_{C})/db$ from AMD calculation. The dashed lines are fits to the distribution with the equation $2\pi b/(1+\exp((b-b_{0})/\Delta b))$, see Ref. [15].
@@ IV Results and discussion
The implementation of in-medium effects is not universal across reactions at different beam energies. In our calculation, $\eta=0.35$ is consistent with data in the reaction at $56$ $\mathrm{MeV}/\mathrm{nucleon}$ and $\eta=0.85$ is consistent with data in the reaction at $140$ $\mathrm{MeV}/\mathrm{nucleon}$, indicating more suppression in the medium in the reaction at $56$ $\mathrm{MeV}/\mathrm{nucleon}$. It implies that the in-medium effect depends not only on the local density but also reflects the dynamic situation of the medium, such as a nontrivial phase space distribution. However, previous study in $\mathrm{Sn}+\mathrm{Sn}$ reactions at $270$ $\mathrm{MeV}/\mathrm{nucleon}$ showed that the $\mathrm{AMD}$ calculation cannot simultaneously reproduce the rapidity distributions and the transverse momentum spectra of light charged particles [19, 22] by only adjusting parameters of the NN collision matrix element.